-X+y=-2
x+y=2
solve for linear equation

Answer by jim_thompson5910(35256)

Add the equations

-x+y=-2
x+y=2
-------
0x+2y=0

2y = 0 ---> y = 0

Use this to find x

-x+y=-2

-x+0=-2

-x = -2

-x = 2

x = 2

The solution is (x,y) = (2,0)